The hypothenuse of a right triangle measures 25 cms. A leg of the triangle measures 15 cms. What is the area of the triangle? please show work/explaition thank you so much :D
--sandy

Answer by checkley71(8403)

25^2=15^2+x^2 or 625=225+x^2 or x^2=625-225 or x^2=400 opr x=20
area of right triangle is A=1/2*b*h or A=1/2*20*15 OR 1/2*300 OR A=150